OpenStudy (anonymous):

Use the discriminant to determine the nature of the roots of 3x2 = 0.

OpenStudy (anonymous):

A:no real roots C:one real root B:two distinct real roots D:three distinct real roots

OpenStudy (anonymous):

@Limitless Do you know how to do this?

OpenStudy (anonymous):

Your discriminant is \(D=b^2-4ac.\) The \(b\), \(a\), and \(c\) come from the equation \(ax^2+bx+c=0.\) Since you have \(3x^2=0,\) you just match the numbers and put them into \(b^2-4ac.\) (Since there isn't a \(bx\) and \(c\), \(b\) and \(c\) are both zero.) Now, what your \(D\) is tells you about the roots. If \(D=0\), there are two real double roots. If \(D>0\), there are two real roots which are not doubles. If \(D<0\), there are two imaginary roots.